Solve the equation (3y+4)(y-4)=(y+20)(3y-20)
Answer by mananth(16946)

(3y+4)(y-4)=(y+20)(3y-20)
3y^2-12y+4y-16=3y^2-20y+60y-400
3y^2-3y^2-12y+4y+20y-60y-16+400=0
-48y=-384
y=-384/-48
y=8